At a local college, 10 of the male students are smokers and 490 are non-smokers. Of the female students, 40 are smokers and 363 are non-smokers. A male student and a female student from the college are randomly selected for a survey. What is the probability that both are smokers?
Do not round your answer.

Answer by htmentor(789)

The fraction of males who are smokers = 10/500 = 1/50
The fraction of females who are smokers = 40/403
The probability that both are smokers is the product of these two:
P(both) = (1/50)*(40/403) = (4/5)*(1/403)